factorise x^3 +2x^2 +2

Grade:10

2 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
Given expression is x^3 -2x^2 + 2
It can also be written as
=x3+2x2 + 2
=x2(x+2)+2
This expression is unfactorizable further and ends here.
SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find my response for your question below
Given =x3 +2x2+2
=x2(x+2)+2
It is not factorable further.
I think you have missed an ‘x’ term in the question that is why it is not getting factored further.please check the question and post it again correctly so that i can provide you with a meaningful answer.